A crane lifts 1, 800 kg load os scrap metal to a distance of 12 m in 20 s. How much power is delivered to the road?


1
Expert's answer
2021-10-06T08:14:00-0400
"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{1800\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot12\ m}{20\ s}=10584\ W=10.6\ kW."